OpenStudy (anonymous):

PLEASE HELP!!! He paddled for 4hrs with 4km/h current. On way back it took 8 hrs against current. Find speed of boat in still water.

OpenStudy (anonymous):

Let the distance be X. Then (using distance = speed times time)) X = (u+4)4. On the way back, he will also travel a distance of X, but will be fighting the current so X=(u-4)8. Solve the equation for u: (u+4)4=(u-4)8 --> -4u = -16-32 --> u = 12.

OpenStudy (mertsj):